OpenStudy (anonymous):

In triangle ABC, BC = 8, CA = 10, and AB = 12. If M is the midpoint of CA, and BP bisects angle B, find MP.

OpenStudy (anonymous):

is therea picturre with this? this is a simple triangle but clarity needs to be on where the points are located,

OpenStudy (anonymous):

No, that its the only information given

OpenStudy (anonymous):

that is*

OpenStudy (anonymous):

i think it is 3 - this the notion that P is located inside the triangle

OpenStudy (anonymous):

How did you get your answer? Can you please explain?

OpenStudy (anonymous):

what level math is this so I can know what you should have been taught thus far

OpenStudy (anonymous):

high school geometry, learning about ratio and proportions and parts of similar triangles

OpenStudy (anonymous):

are you even there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

What have you been doing all this time? lol

OpenStudy (anonymous):

i am thinking of how to do this

OpenStudy (anonymous):

I don't think the picture right about point P

OpenStudy (anonymous):

Bisect angle does NOT divide opposite side evenly!

OpenStudy (accessdenied):

i have this uncanny urge to use coordinate geometry to place this triangle on a plane and find equations until something relevant happens

OpenStudy (anonymous):

You're way too .. hold on, I'm busy

OpenStudy (anonymous):

I'm back! I'm thinking of applying the golden ratio rule because it's not right triangle!

OpenStudy (accessdenied):

M(16.5, sqrt(43.25)/2) this seems like a lot more than i would expect in geometry perhaps it was not meant to be... golden ratio rule? this is news to me

OpenStudy (anonymous):

It's a rule applies for bisect angle!

OpenStudy (anonymous):

Very interesting, I learn it from solving recent several geometry problems

OpenStudy (anonymous):

a/b = (a + b)/a

OpenStudy (anonymous):

You should check out the bisect angle to see illustrated picture!

OpenStudy (anonymous):

Since AC is divided into half MC and MA, then the bisect angle line divides MC into PC and PM

OpenStudy (anonymous):

In other thought: 8/ x = 12/ ( 10 -x) 12x = 80 - 8x ->20x = 80 => x = 4 ( AP) then 10 - x = 6 ( BP) Thus PM = PB - BM = 6 - 5 = 1

OpenStudy (anonymous):

@AccessDenied check it out!